# Recovering Injected Astrophysics from the LISA Galactic Binaries

### v1

@@ 1 Introduction
In the past decade, the LIGO/Virgo/KAGRA collaboration has published the first observations of gravitational waves from compact binary mergers in the kilo-Hertz frequency band (Abbott et al., 2016). These measurements probe populations of stellar mass binary black holes and neutron stars which electromagnetic observations cannot probe except by targeting X-ray emission from accreting binaries (Fabbiano, 2006; Remillard & McClintock, 2006; Miller-Jones, 2014; Inoue et al., 2022; Fortin, Francis et al., 2023). However, X-ray emission catches only a subset of all the compact object binaries, particularly those that are actively accreting at the current time. In the 2030s, the Laser Interferometer Space Antenna (LISA) will be the first space-based gravitational wave interferometer (eLISA Consortium et al., 2013). LISA will be most sensitive in a lower gravitational wave frequency than its extant ground-based counterparts and will observe detached compact binaries, such as the Double White Dwarf (DWD) population in The Milky Way (Nelemans et al., 2001; Ruiter et al., 2010; Toonen et al., 2012; Korol et al., 2017, 2018; Lamberts et al., 2019; Breivik et al., 2020; Korol et al., 2022b) .
@@
The LISA Galactic DWD binaries will probe isolated binary evolution in the lower end of the stellar mass spectrum where most stars reside. Several million DWDs in the Milky Way and local group are expected to radiate gravitational waves in the LISA band, with the majority having overlapping frequencies that can be observed as a Galactic gravitational-wave foreground (e.g. Benacquista & Holley-Bockelmann, 2006; Breivik et al., 2020; Staelens & Nelemans, 2024; Hofman & Nelemans, 2024). In addition to the Galactic foreground, is expected that LISA will discover tens-of-thousands of individually resolved DWDs in the Milky Way (e.g. Nelemans et al., 2001; Nissanke et al., 2012; Tang et al., 2024) and the local group (e.g. Korol et al., 2018; Roebber et al., 2020; Korol et al., 2020; Keim et al., 2023). In addition to isolated binaries, LISA will discover DWDs that have been formed either dynamically in environments like globular clusters (Willems et al., 2007; Kremer et al., 2018), or higher-multiplicity systems in the field of the Milky Way (e.g. Offner et al., 2023). However, the Type Ia supernova rate from merging DWDs in globular clusters may be as low as $10^{-7}yr^{-1}$ in an environment like the Milky-Way (Kremer et al., 2020, 2021), compared to a $10^{-3}yr^{-1}$ estimate for the Type Ia rates of Milky-Way-like galaxies (Frohmaier et al., 2018). Furthermore, the triple fraction of WD progenitor stars is measured to make up $20\%$ or less of the stellar population (e.g. Offner et al., 2023). Thus, the majority of resolved DWDs discovered by LISA are expected to have been born in isolated binary systems. Because of this, we can use comparisons between the predicted and observed isolated binary populations to infer constraints on models of binary evolution and learn about the underlying astrophysical processes.
@@
As proof of concept, comparable methods have been applied to interpret the population of sources identified by ground-based gravitational wave detectors. The LVK has published approximately ninety gravitational wave observations in the Gravitational Wave Transient Catalog (GWTC) (Abbott et al., 2019, 2021a; The LIGO Scientific Collaboration et al., 2021a; The LIGO Scientific Collaboration & The Virgo Collaboration, 2021). The population of these observations informs us about the underlying population of merging compact binaries (The LIGO Scientific Collaboration et al., 2020, 2021b; Abbott et al., 2021b). These merger events are expected to form from specific formation channels, such as the isolated evolution of massive stellar binaries (Hurley et al., 2002; Eldridge et al., 2017; Giacobbo et al., 2017; Breivik et al., 2020; Belczynski et al., 2020; Broekgaarden & Berger, 2021; Broekgaarden et al., 2021; Fragos et al., 2022; Stevenson & Clarke, 2022; Delfavero et al., 2023) , and various dynamic formation scenarios (Romero-Shaw et al., 2020; Fragione et al., 2020; Gayathri et al., 2022; Gamba et al., 2023); (see Mandel & Farmer, 2022) for a detailed review of formation channels. By comparing gravitational wave observations to the results of binary evolution population synthesis models, we can use Bayesian inference to learn about the underlying astrophysics (Zevin et al., 2021; Stevenson et al., 2017; Delfavero et al., 2023). Within the context of these ground-based instruments, previous studies have used synthetic populations of sources within an end-to-end calculation to assess how well these calculations can constrain source model parameters, a process we denote for brevity as “injection recovery” (see, e.g., Barrett et al., 2017; Wysocki et al., 2019; Zevin et al., 2021) and references therein.
@@
2 Simulating a population of galactic binaries
There are many uncertainties remaining in the evolution of stellar binaries, such as the initial binary fraction (Thiele et al., 2023) and various forms of mass transfer initiated by Roche-overflow interactions. One such form of dynamically unstable mass transfer is common envelope evolution, where the atmosphere of a donor star expands beyond its Roche-lobe such that both stars become immersed in a common envelope. The most accurate method for characterizing the efficiency of common envelope evolution (in ejecting orbital energy and angular momentum from the system) is currently debated by various groups ((e.g. Ivanova, 2011; Ivanova & Chaichenets, 2011; Klencki et al., 2021; Wilson & Nordhaus, 2022)).
@@ 2.1 Binary Population Synthesis with COSMIC
The primary mass of each binary is sampled from a Kroupa 2001 IMF (Kroupa & Weidner, 2003), while the secondary is drawn from a uniform distribution in mass ratio ($q\in[0.01,1.0]$). COSMIC assumes a metallicity-dependent close binary fraction consistent with Moe et al. (2019) that decreases from $50\%$ at Z=0.0001 to $15\%$ at Z=0.02. The orbital periods are distributed log-normally with increased numbers of systems below $10,000\,\rm{days}$ at sub-solar metallicities. We further assume that initial orbital eccentricities are uniformly distributed between 0 and 1.
@@ 2.2 Gravitational wave evolution and delay time
Ultimately, the quantities which describe potential gravitational wave observations are masses and orbital periods. These are related to the star forming conditions via a delay time: the time between the zero age main sequence formation of the binary and a gravitational wave observation. To track this evolution properly, we characterize each binary by the separation and mass of each remnant at the time of the formation of the white dwarf. As DWD orbital frequencies continue to evolve over time according to gravitational wave radiation, we calculate this “delay time” as the sum of the time required to form the compact binary ($t_{\mathrm{form}}$) and the time required to evolve the system to a final orbital frequency at present day ($t_{\mathrm{evol}}$). Therefore, we draw several quantities from the output of COSMIC simulations: the separation of the binary at DWD formation, the mass of each remnant, and the time of compact binary formation. Because Roche-lobe overflow is assumed to fully circularize close DWDs, which are the products of either stable mass transfer or common envelope, all DWD binaries analyzed in this work are assumed to form as circular binaries. We also interpret COSMIC outputs to track the prevalence of different chemical compositions in the underlying white dwarf populations.
@@
Gravitational wave radiation can drive some binaries to transfer mass or contact as a DWD binary before present day. For these systems, the “delay time” is the time between star formation and contact (rather than between star formation and present day). The fate of such a system can be hard to determine given our current state of knowledge, but may depend on many things including the mass ratio of the binary (Marsh et al., 2004; Gokhale et al., 2007; Sepinsky et al., 2010; Kremer et al., 2015, 2017). While these binaries may be observable by LISA, they will be identified separately from the detached binary population (Breivik et al., 2018), and we consider them beyond the scope of this work. Our assumptions regarding the critical separation of the DWD binaries before mass transfer may occur are consistent with Thiele et al. (2023).
@@ 2.3 Results of changing $\alpha_{\mathrm{CE}}$ on binary formation pathways
Figure 1 exhibits the formation efficiency of DWD binaries for a few choices of $\alpha_{\mathrm{CE}}$, selected to encompass the range of all simulated alphas. We note that as $\alpha_{\mathrm{CE}}$ increases, $\eta_{\mathrm{form}}$ increases as well; broadly, this is because fewer systems undergo pre-DWD stellar mergers when common envelope ejection is more efficient. We point the reader to Section 5 of Thiele et al. (2023) for a related discussion of formation efficiency trends. We also note that this scaling is not applied uniformly across all kinds of DWD binaries and metallicities (note the varied y-axis ranges in Figure 1).. The common envelope ejection efficiency has a stronger impact on binaries containing He WDs becuase their progenitors have smaller radii relative to CO or ONe WD progenitors. This means that the common envelopes which produce He WDs occur in orbits with initially shorter periods and thus a smaller orbital energy well to be used in unbinding the envelope from the system. As the common envelope ejection efficiency decreases, we find a significant reduction in the number of DWDs containing He WD due to mergers caused by  unsuccessful envelope ejection. We can therefore search for a dependence on $\alpha_{\mathrm{CE}}$ in both the total number of binaries and the relavence of different modes in their orbital parameters (or the shape of the density).
@@ 3 A synthetic galaxy of LISA sources
In order to simulate a LISA population of DWD binaries, we need to sample from star formation assumptions in a synthetic galaxy. We fill this synthetic galaxy with binaries simulated by COSMIC to construct a galactic DWD population. Following this, we classify binaries based on whether they have time to form and remain in the LISA band until present day. Finally, the binaries resolved above the confusion limit are identified from the sample population and their uncertainties are estimated using a Fisher matrix method with the ldasoft code (Littenberg et al., 2020).
@@ 3.1 Sampling star formation with FIRE-2 m12i
Within COSMIC, solar metallicity ($\mathcal{Z}_{\odot}=0.02$) is consistent with Tout et al. (1996), to which the the radii and luminosity power-laws from BSE are calibrated. This Tout et al. (1996) metallicity assumption is a poor estimator of iron abundances, as elemental abundances are estimated from meteorites (Anders & Grevesse, 1989). Meanwhile, in FIRE-2, solar metallicity is assumed to be $\mathcal{Z}_{\odot}=0.01342$, calibrated to Asplund et al. (2009). For this reason, absolute metallicity is used to inform this metallicity-dependent sampling process, as using different solar metallicity assumptions for FIRE-2 and COSMIC star formation can otherwise introduce a bias.
@@ 3.2 Confusion Limited Sources
We use this Gaussian approximation to assume the form of our likelihood function $\mathcal{L}(\vec{\lambda})$, where our binary parameters are gravitational wave frequency ($f_{\mathrm{GW}}$), the time derivative of that frequency ($\dot{f}_{\mathrm{GW}}$), and strain amplitude ($h$). Following this, we use the Jacobian transformation to change coordinates to $f_{\mathrm{GW}}$, chirp mass ($\mathcal{M}_{c}$), and distance ($\mathrm{\ell}$). We then marginalize over distance to construct a likelihood for each source, $\mathcal{L}(f_{\mathrm{gw}},\mathcal{M}_{c})$. We acknowledge the limitations of Fisher estimates for gravitational wave parameter uncertainties (Vallisneri, 2008). However, as the number of sources can vary widely (up to tens of millions of sources), estimating a detailed likelihood approximation more detailed than a Fisher estimate for each source is beyond the scope of this work.
@@ 4.1 The Inhomogeneous Poisson Point Process likelihood
$\mathrm{P}(\mathcal{H}|d)=\frac{\mathrm{P}(d|\mathcal{H})\mathrm{P}(\mathcal{H})}{\mathrm{P}(d)},$ (7)
. where $\mathrm{P}(\mathcal{H}|d)$ is the posterior probability of our hypotheses, $\mathrm{P}(d|\mathcal{H})$ is the likelihood of our data given our set of hypotheses, $\mathrm{P}(\mathcal{H})$ is the probabilistic representation of our prior knowledge about $\mathcal{H}$, and $\mathrm{P}(d)$ is a marginalization factor to keep our probability distribution normalized. That marginalization factor can be calculated over all possible hypotheses:
@@
The Inhomogeneous Poisson Point Process extends the treatment of a counting experiment, describing the likelihood for both the event rate, and the properties of a set of observations, compared to a given model (Moller & Waagepetersen, 2003). This is especially useful when our observations ($\{d_{j}\}$) have parameter uncertainties characterized by a source likelihood ($\mathcal{L}_{j}(\vec{\lambda})=\mathrm{P}(d_{j}|\vec{\lambda},\Lambda)$). The Inhomogeneous Poisson Point Process likelihood is the probability, $\mathrm{P}(\{d_{j}\}|\Lambda)$, that the set of individual detections $d_{j}$ could be observed under an assumed population model characterized by some assumptions $\Lambda$.

### v2

# Recovering Injected Astrophysics from the LISA Double White Dwarf Binaries

@@ 1 Introduction
In the past decade, the LIGO/Virgo/KAGRA collaboration has published the first observations of gravitational waves from compact binary mergers in the kilo-Hertz frequency band (Abbott et al., 2016). These measurements probe populations of stellar mass binary black holes and neutron stars which electromagnetic observations cannot probe except by targeting X-ray emission from accreting binaries (Fabbiano, 2006; Remillard & McClintock, 2006; Miller-Jones, 2014; Inoue et al., 2022; Fortin, Francis et al., 2023). In the 2030s, the Laser Interferometer Space Antenna (LISA) will be the first space-based gravitational wave interferometer (eLISA Consortium et al., 2013). LISA will be most sensitive in a lower gravitational wave frequency than its extant ground-based counterparts and will observe detached compact binaries, such as the Double White Dwarf (DWD) population in The Milky Way (Nelemans et al., 2001; Ruiter et al., 2010; Toonen et al., 2012; Korol et al., 2017, 2018; Lamberts et al., 2019; Breivik et al., 2020; Korol et al., 2022b; Hirai & Mandel, 2022; Biscoveanu et al., 2023; Tang et al., 2024) .
@@
The LISA Galactic DWD binaries will probe isolated binary evolution in the lower end of the stellar mass spectrum where most stars reside. Several million DWDs in The Milky Way and local group are expected to radiate gravitational waves in the LISA band, with the majority having overlapping frequencies that can be observed as a Galactic gravitational-wave foreground (e.g. Benacquista & Holley-Bockelmann, 2006; Breivik et al., 2020; Staelens & Nelemans, 2024; Hofman & Nelemans, 2024). In addition to the Galactic foreground, is expected that LISA will discover tens-of-thousands of individually resolved DWDs in The Milky Way (e.g. Nelemans et al., 2001; Nissanke et al., 2012; Tang et al., 2024) and the local group (e.g. Korol et al., 2018; Roebber et al., 2020; Korol et al., 2020; Keim et al., 2023). In addition to isolated binaries, LISA will discover DWDs that have been formed either dynamically in environments like globular clusters (Willems et al., 2007; Kremer et al., 2018), or higher-multiplicity systems in the field of The Milky Way (e.g. Offner et al., 2023). However, the Type Ia supernova rate from merging DWDs in globular clusters may be as low as $10^{-7}yr^{-1}$ in an environment like The Milky Way (Kremer et al., 2020, 2021), compared to a $10^{-3}yr^{-1}$ estimate for the Type Ia rates of Milky-Way-like galaxies (Frohmaier et al., 2018). Furthermore, the fraction of WD progenitor stars formed in stellar triples is measured to make up $20\%$ or less of the stellar population (e.g. Offner et al., 2023). Thus, the majority of resolved DWDs discovered by LISA are expected to have been born in isolated binary systems. Because of this, we can use comparisons between the predicted and observed isolated binary populations to infer constraints on models of binary evolution and learn about the underlying astrophysical processes.
@@
As proof of concept, comparable methods have been applied to interpret the population of sources identified by ground-based gravitational wave detectors. The LVK has published approximately ninety gravitational wave observations in the Gravitational Wave Transient Catalog (GWTC) (Abbott et al., 2019, 2021a; The LIGO Scientific Collaboration et al., 2021a; The LIGO Scientific Collaboration & The Virgo Collaboration, 2021). The population of these observations informs us about the underlying population of merging compact binaries (The LIGO Scientific Collaboration et al., 2020, 2021b; Abbott et al., 2021b). These merger events are expected to form from specific formation channels, such as the isolated evolution of massive stellar binaries (Hurley et al., 2002; Eldridge et al., 2017; Giacobbo et al., 2017; Breivik et al., 2020; Belczynski et al., 2020; Broekgaarden & Berger, 2021; Broekgaarden et al., 2021; Fragos et al., 2022; Stevenson & Clarke, 2022; Delfavero et al., 2023) , and various dynamic formation scenarios (Romero-Shaw et al., 2020; Fragione et al., 2020; Gayathri et al., 2022; Gamba et al., 2023); see Mandel & Farmer (2022) for a detailed review of formation channels. By comparing gravitational wave observations to the results of binary evolution population synthesis models, we can use Bayesian inference to learn about the underlying astrophysics (Zevin et al., 2021; Stevenson et al., 2017; Delfavero et al., 2023). Within the context of these ground-based instruments, previous studies have used synthetic populations of sources within an end-to-end calculation to assess how well these calculations can constrain source model parameters, a process we denote for brevity as “injection recovery” (see, e.g., Barrett et al., 2017; Wysocki et al., 2019; Zevin et al., 2021) and references therein.
@@
2 Simulating a population of Galactic DWD binaries
There are many uncertainties remaining in the evolution of stellar binaries, such as the initial binary fraction (Thiele et al., 2023) and various forms of mass transfer initiated by Roche overflow interactions. One such form of dynamically unstable mass transfer is common envelope evolution, where the atmosphere of a donor star expands beyond its Roche lobe such that both stars become immersed in a common envelope. The most accurate method for characterizing the efficiency of common envelope evolution (in ejecting orbital energy and angular momentum from the system) is currently debated by various groups (e.g. Ivanova, 2011; Ivanova & Chaichenets, 2011; Klencki et al., 2021; Wilson & Nordhaus, 2022; Hirai & Mandel, 2022; Stefano et al., 2023; Tang et al., 2024). If the common envelope prescription employed in the simulation of DWD binaries fails to describe the onset and outcomes of common envelope evolution, this may account for a mismatch between the properties of observed and simulated populations.
@@ 2.1 Binary Population Synthesis with COSMIC
The primary mass of each binary is sampled from a Kroupa 2001 IMF (Kroupa & Weidner, 2003), while the secondary is drawn from a uniform distribution in mass ratio ($q\in[0.01,1.0]$). COSMIC assumes a metallicity-dependent close binary fraction (i.e. fraction of stellar binaries with initial orbital periods less than six days) consistent with Moe et al. (2019) that decreases from $50\%$ at Z=0.0001 to $15\%$ at Z=0.02. The orbital periods are distributed log-normally with increased numbers of systems below $10,000\,\rm{days}$ at sub-solar metallicities. We further assume that initial orbital eccentricities are uniformly distributed between 0 and 1.
@@ 2.2 Gravitational wave evolution and delay time
Ultimately, the quantities which describe potential gravitational wave observations are masses and orbital periods. These are related to the star forming conditions via a delay time: the time between the zero age main sequence formation of the binary and a gravitational wave observation. To track this evolution properly, we characterize each binary by the separation and mass of each remnant at the time of the formation of the white dwarf. As DWD orbital frequencies continue to evolve over time according to gravitational wave radiation, we calculate this “delay time” as the sum of the time required to form the compact binary ($t_{\mathrm{form}}$) and the time required to evolve the system to a final orbital frequency at present day ($t_{\mathrm{evol}}$). Therefore, we draw several quantities from the output of COSMIC simulations: the separation of the binary at DWD formation, the mass of each remnant, and the time of compact binary formation. Because Roche lobe overflow is assumed to fully circularize close DWDs, which are the products of either stable mass transfer or common envelope, all DWD binaries analyzed in this work are assumed to form as circular binaries. We also interpret COSMIC outputs to track the prevalence of different chemical compositions in the underlying white dwarf populations.
Gravitational wave radiation can drive some binaries to transfer mass or contact before present day. These systems can exchange mass or merge during the DWD stage of a binary’s life. For these systems, the “delay time” is the time between star formation and contact (rather than between star formation and present day). The fate of such a system can be hard to determine given our current state of knowledge, but may depend on many things including the mass ratio of the binary (Marsh et al., 2004; Gokhale et al., 2007; Sepinsky et al., 2010; Kremer et al., 2015, 2017). While these binaries may be observable by LISA, they will be identified separately from the detached binary population (Breivik et al., 2018), and we consider them beyond the scope of this work. Our assumptions about how close a DWD binary must come (i.e. the critical separation) before mass transfer may occur are consistent with Thiele et al. (2023).
@@ 2.3 Results of changing $\alpha_{\mathrm{CE}}$ on binary formation pathways
Figure 1 exhibits the formation efficiency of DWD binaries for a few choices of $\alpha_{\mathrm{CE}}$, selected to encompass the range of all simulated alphas. We note that as $\alpha_{\mathrm{CE}}$ increases, $\eta_{\mathrm{form}}$ increases as well; broadly, this is because fewer systems undergo pre-DWD stellar mergers when common envelope ejection is more efficient. We point the reader to Section 5 of Thiele et al. (2023) for a related discussion of formation efficiency trends. We also note that this scaling is not applied uniformly across all kinds of DWD binaries and metallicities (note the varied y-axis ranges in Figure 1). The common envelope ejection efficiency has a stronger impact on binaries containing He WDs because their progenitors have smaller radii relative to CO or ONe WD progenitors. This means that the common envelopes which produce He WDs occur in orbits with initially shorter periods and thus a smaller orbital energy well to be used in unbinding the envelope from the system. As the common envelope ejection efficiency decreases, we find a significant reduction in the number of DWDs containing He WD due to mergers caused by  unsuccessful envelope ejection. This motivates a search for further quantitative (statistical) dependence on $\alpha_{\mathrm{CE}}$ through i) the total number of binaries, ii) the relative contribution of different formation pathways to the overall population, and iii) the shape of the density in observable quantities.
@@ 3 A synthetic galaxy of LISA sources
In order to simulate a LISA population of DWD binaries, we need to assume a star formation history by sampling from star formation in a synthetic galaxy. We fill this synthetic galaxy with binaries simulated by COSMIC to construct a Galactic DWD population. Following this, we classify binaries based on whether they have time to form and remain in the LISA band until present day. Finally, the binaries resolved above the confusion limit are identified from the sample population and their uncertainties are estimated using a Fisher matrix method with the ldasoft code (Littenberg et al., 2020).
@@ 3.1 Sampling star formation with FIRE-2 m12i
Within COSMIC, solar metallicity ($\mathcal{Z}_{\odot}=0.02$) is consistent with Tout et al. (1996), to which the the radii and luminosity power-laws from BSE are calibrated. This Tout et al. (1996) metallicity assumption is a poor estimator of iron abundances, as elemental abundances are estimated from meteorites (Anders & Grevesse, 1989). Meanwhile, in FIRE-2, solar metallicity is assumed to be $\mathcal{Z}_{\odot}=0.01342$, calibrated to Asplund et al. (2009). For this reason, absolute metallicity (i.e. the base 10 logarithm of the fraction of elements with an atomic number of 3 or higher) of m12i star particles is used to inform this metallicity-dependent sampling process, as using different solar metallicity assumptions for FIRE-2 and COSMIC star formation can otherwise introduce a bias.
@@ 3.2 Confusion Limited Sources
We use this Gaussian approximation to assume the form of our likelihood function $\mathcal{L}(\vec{\lambda})$, where our binary parameters are gravitational wave frequency ($f_{\mathrm{GW}}$), the time derivative of that frequency ($\dot{f}_{\mathrm{GW}}$), and strain amplitude ($h$). Following this, we use the Jacobian transformation to change coordinates to $f_{\mathrm{GW}}$, chirp mass ($\mathcal{M}_{c}$), and distance ($\mathrm{\ell}$). We then marginalize over distance to construct a likelihood for each source, $\mathcal{L}(f_{\mathrm{gw}},\mathcal{M}_{c})$. In gravitational wave Parameter Estimation (PE), Fisher estimates for parameter uncertainties can encounter numerical challenges, as well as bias (when astrophysical prior information is not fully considered). These and other limitations of Fisher estimates are summarized well by (Vallisneri, 2008). However, as the number of sources can vary widely (up to tens of millions of sources), estimating a detailed likelihood approximation more detailed than a Fisher estimate for each source is beyond the scope of this work.
@@ 4.1 The Inhomogeneous Poisson Point Process likelihood
$\mathrm{P}(\mathcal{H}|d)=\frac{\mathrm{P}(d|\mathcal{H})\mathrm{P}(\mathcal{H})}{\mathrm{P}(d)}.$ (7)
where $\mathrm{P}(\mathcal{H}|d)$ is the posterior probability of our hypotheses, $\mathrm{P}(d|\mathcal{H})$ is the likelihood of our data given our set of hypotheses, $\mathrm{P}(\mathcal{H})$ is the probabilistic representation of our prior knowledge about $\mathcal{H}$, and $\mathrm{P}(d)$ is a marginalization factor to keep our probability distribution normalized. That marginalization factor can be calculated over all possible hypotheses:
@@
The Inhomogeneous Poisson Point Process extends the treatment of a counting experiment, describing the likelihood for both the event rate, and the properties of a set of observations, compared to a given model (Moller & Waagepetersen, 2003). This is especially useful when our observations ($\{d_{j}\}$) have parameter uncertainties characterized by a source likelihood:
$\mathcal{L}_{j}(\vec{\lambda})=\mathrm{P}(d_{j}|\vec{\lambda},\Lambda)\;.$ (12)
Let the indices $j$ and $k$ refer to a sample binary from an observed (or injected) and modeled population respectively. The Inhomogeneous Poisson Point Process likelihood is the probability, $\mathrm{P}(\{d_{j}\}|\Lambda)$, that the set of individual detections $d_{j}$ could be observed under an assumed population model characterized by some assumptions $\Lambda$.